# SymbolicAI: A framework for logic-based approaches combining generative models and solvers

### v2

@@ 3 Problem Definition
Conventional approaches employing foundation models for inference, such as LLMs, are predominantly confined to single-step or few-step executions and primarily reliant on hand-crafted in-context learning prompt instructions. This restricted scope limits the utilization to single modalities, lacks refinement or verification, and exhibits limited tool proficiency. We posit that the integration of NeSy engines as core computation units, realized through logic-based methodologies coupled with sub-symbolic foundation models, offers a more general, robust, and verifiable perspective. This approach has several advantages. Firstly, it facilitates the integration of pre-existing engineered solutions (e.g. various classical algorithms), offloading computational complexity and bridging various modalities. Secondly, it enables sub-symbolic generalization to focus on evidence-based decision-making (e.g. selecting the respective tool based on in-context classification). Thirdly, it provides an interpretable language-based control layer for explainable, autonomous systems. Central to our solution is a method to define and measure the orchestration of interactions between symbolic and sub-symbolic systems, and the level at which instructions are formulated for effective control and task execution.
@@ 4 Design Principles
In the following subsections, we elaborate on the key design principles underlying SymbolicAI and how we guide the generative processes of NeSy engines. For further technical details, see Section 5.
@@
As posited by Newell & Simon (1976), symbols are elemental carriers of meaning within a computational context333 We base our framework’s name on the aspirational work of Newell and Simon.. These symbols define physical patterns capable of composing complex structures, and are central to the design and interpretation of logic and knowledge representations (Augusto, 2022). Thus, SymbolicAI conceptualizes the notion that symbols, and the expressions they form, are reflections of the information inherent in a system, and serve as surrogate for the interaction between the system and the problem space. Moreover, we argue that real patterns, as Dennett (1991) speaks of, can be effectively realized through the use of symbols because these symbols act as versatile abstractions that capture and represent the underlying structures and dynamics of these patterns, facilitating their interpretation and manipulation in computational models.
Furthermore, we attribute task-specific mappings to a language-centric strategy, leveraging their inherent semantics and abstraction to describe the states and properties of the problem at hand. These mappings are universal and may be used to define scene descriptions, long-horizon planning, acoustic properties, emotional states, physical conditions, etc. Therefore, we adhere to the analogy of language representing the convex hull of the knowledge of our society, utilizing it as a fundamental tool to define symbols. This approach allows us to map the complexities of the world onto language, where language itself serves as a comprehensive, yet abstract, framework encapsulating the diversity of these symbols and their meanings. This perspective resonates with our innate human tendencies to attribute existing physical objects with abstract concepts, as exemplified by our natural inclination to link tangible objects to colors and emotions, such as blending the color ”red” with ”heart”, ”warm”, and ”passion”.
However, this language-centric model does not inherently encompass all forms of representation, such as sensory inputs and non-discrete elements, requiring the establishment of additional mappings to fully capture the breadth of the world. This limitation is manageable, since we care to engage in operations within this abstract conceptual space, and then define corresponding mappings back to the original problem space. These are typically applied through function approximation, as in typical modality-to-language and language-to-modality use cases, where modality is a placeholder for various skill sets such as text, image, video, audio, motion, etc.
Ultimately, this approach also anchors our work in the field of formal language theory, as we require a structured method to construct mappings from the world to language. This grounding sets the foundation for employing formal language structures, such as grammars, to systematically define our language-centric approach to problem-solving and the associated translation of real-world complexities into linguistic terms.
@@
In formal language theory and linguistics, languages are structured following the Chomsky hierarchy, which classifies languages by the complexity of their grammatical structure (Chomsky, 1956). This hierarchy, comprising four types of grammars (Type-3 to Type-0), delineates formal languages by their grammatical complexity. A grammar in this context consists of terminal and non-terminal symbols, production rules, and a designated start symbol, facilitating the generation of valid strings within a language. In developing SymbolicAI, we posit that all symbols can be represented as strings, augmented with conditional instructions and types derived from a domain-specific language (DSL) tailored for directing NeSy computation engines, like LLMs (see Figure 2).
A key advancement of LLMs over previous systems lies in their ability to generalize from formal languages (Wang et al., 2023a) and knowledge systems, primarily due to their world-knowledge and proficiency in understanding context-based analogies. While there is currently no universal consensus among experts regarding the precise classification of natural language within the Chomsky hierarchy, we have seen in our empirical evaluations promising results utilizing LLMs as semantic parsers. This approach can be viewed as employing a form of flexible, context-sensitive grammar, which enables the processing of instructions and analogies with a nuanced understanding of language’s inherent variability and complexity. The intersection between formal and natural languages becomes evident when considering how language patterns, through prompts like ”You are a helpful assistant…”, elicit structured responses, indicating a potential underlying formal mechanism at play. This observation underlines the utility of such a grammar in our framework, particularly within in-context learning, where it serves as an explicit schema guiding the structure of examples used in few-shot learning scenarios. For instance, equating ”3.1415…” with ”$\pi$” or ”August 4, 1961” with ”1961-08-04” in a given context demonstrates this probabilistic, context-dependent interpretation of symbols. Such a system doesn’t rigidly adhere to standard grammatical rules but instead adjusts and interprets based on the presented context, effectively creating a dynamic and situation-specific grammar.
In-Context Learning
Recently, several in-context learning methodologies evolved to enable tool usage through LLMs (Schick et al., 2023), or refine the generative outcome of LLMs (Yang et al., 2023). This includes chain-of-thought (CoT) prompting, a method that conditions the model to reveal its step-by-step reasoning process (Wei et al., 2022b; Singhal et al., 2023). CoT prompting breaks down complex tasks into simpler, sequential steps, and helps with interpreting LLM’s output. Self-generated CoT, where models are encouraged to generate their own reasoning chains based on training examples, surpasses even expertly crafted CoT (Fernando et al., 2023). This observation echoes other reports that GPT-4 has an emergent self-improving capability through introspection, such as self-verification (Weng et al., 2023) or self-consistency (Wang et al., 2023b). Tree of Thoughts (ToT) enables LLMs to solve complex problems by exploring multiple reasoning paths through a search tree of coherent text units, demonstrating significant problem-solving enhancements in tasks requiring strategic planning and search (Yao et al., 2023a). Ensemble techniques further enhance the robustness and accuracy of model predictions by combining several strategies to establish a consensus (Nori et al., 2023). Conceptually, we designed our framework to enable all these techniques and combine them into dedicated components and sub-processes. Figure 2 outlines a typical expression evaluation in our NeSy pipeline.
@@
Figure 2: Illustration for NeSy pipeline, showcasing conceptual usage of in-context learning methodologies, domain-specific language (DSL) structures, and the expression evaluations through NeSy engines. The expression showcases the left shift operator $\ll$ and how the information of the symbol B is included in the symbol AC. The violet placeholder in the DSL Prompt represents an instruction, such as ”Insert the right-hand side value into the left-hand value in a chronological order.” The positions below represent task-specific few-shot examples.
@@
In-context learning enabled LLMs to become versatile task solvers by interpolating within the training distribution, to the extent that even potentially unseen tasks are addressable (Brown et al., 2020). We attribute this to associations formed within the input space and the capacity of Transformer architectures for defining domain-invariant feature sub-spaces. This phenomenon has striking parallels with advancements in few-shot learning approaches such as SubGD (Gauch et al., 2022), a method based on identifying and utilizing a low-dimensional subspace, learned from various tasks, that effectively acts to regularize the learning process, representing features that are invariant across different learning tasks. Furthermore, SubGD reflects the potential of in-context learning when combined with task-specific fine-tuning by showing that fine-tuning within a learned subspace significantly outperforms traditional fine-tuning methods. We believe that the extent of in-context learning is not yet exhausted, holding considerable promise when used alongside with task-specific fine-tuning and solvers. To develop learning and reasoning systems capable of general problem-solving, we adopt a hybrid methodology. This approach leverages the in-context generalization capability of LLMs, constructing symbolic associations that aim to preserve and propagate situational context, and validating solutions with established solvers.
Function Composition
In SymbolicAI, function composition is relevant for constructing complex hierarchies and behaviors from more basic, fundamental elements. It enables our framework to model interconnected processes, where the output of one function seamlessly transitions into the input of another, thus creating a cohesive sequence of operations. Through function composition, we construct computational graphs, in which intermediate symbols represent the nodes or states within these graphs. Formally, function composition is denoted by $\circ$, where combining functions $f$ and $g$ yields a new function $h=g\circ f$, defined as $h(x)=g(f(x))$ For functions $f:X\rightarrow Y$ and $g:Y\rightarrow Z$, their composition results in a function mapping elements from domain $X$ to codomain $Z$ through $g(f(x))$.
Although traditionally the codomain of the inner function $f$ aligns with the domain of the outer function $g$, SymbolicAI relaxes this constraint by allowing for any subset relationship between these domains and codomains, enhancing data flow flexibility. For example, this relaxed constraint in domain and codomain alignment is particularly beneficial for in-context learning. By leveraging functional few-shot learning, where few-shot examples act as dynamic elements of the function’s domain, SymbolicAI enhances its ability to interpret and respond to diverse input contexts. For instance, a function can classify a user request and select an appropriate engine to process the request. The target modality may vary based on the respective engine. Therefore, in SymbolicAI, the use of function composition leads to the creation of richer and more nuanced symbol-based generative flows, where each functional unit is capable of representing either logical or data-driven behaviors. Importantly, function composition is not confined to strictly symbolic representations; it also connects with sub-symbolic processes. This enables SymbolicAI to handle complex data streams, establishing function composition as a central tenet in bridging multiple modalities and coordinating a variety of tasks.
5 Framework
In this section, we discuss the design choices pertaining to implementation and the specifics of the framework we’ve employed. For an extended overview, see our supplementary materials Section C.
@@ 5 Framework
Analogous to the Python object type, the base type of SymbolicAI is a symbol represented through its name equivalent base type Symbol. A Symbol object marks a non-reducible atomic unit. All other subtypes, such as Expression and its derivatives, are analogous to their mathematical namesakes, representing expressions or units that can be further evaluated and simplified. These subtypes inherit from Symbol the base attributes, primitive operators, and helper methods. Furthermore, each Symbol object contains valued and vector-valued representations, obtained through value and embedding attributes. The latter, in particular, serve as a means to attribute a symbol’s current context, akin to embedding text and storing it as a PyTorch tensor (Paszke et al., 2019) or NumPy array (Harris et al., 2020). While for an LLM, the numerical tensors may lack inherent meaning, vector-valued representations play a strategic role when 1) composite symbols coalesce into more complex expressions, and 2) these embedded tensors become amenable to updates through gradient-based optimization. This design choice is critical in the dynamic landscape of symbolic interactions, particularly for promising applications, such as developing automated and self-evolving NeSy systems.
To enable the processing of symbols by LLMs, we assume that each Symbol object is represented through Python’s native string functionality, where the __str__ method acts an interpretable string representation. Worth noting is that encoding a complex object into a string sometimes precludes the object reconstitution. However, this concern does not substantially impede our methodology: we can employ approximations or proxy representations stored by the vector-valued property to effectively re-map objects. These representations are obtained through respective embedding models. Therefore, we can theoretically assert that any Python object is parsable by an LLM by design. For more details, see our supplementary materials Section E.
@@
Polymorphism is a central concept in programming language theory and prominently featured in SymbolicAI, with significant implications for the design and functionality of our NeSy architecture. Polymorphism refers to the ability of different objects to be accessed through the same interface, or of a single identifier to represent different types based on the context of execution. The provision of a single interface to entities of different types allows operations to be performed in ways specific to their derived types. We employ a polymorphic structure for instruction composition. We designed the Symbol object to contain a global context, which is composed of static and dynamic context parts. The static context is class dependent and defined at design time. The dynamic context is runtime adaptable and can be changed to adhere to runtime specific logic and changes. Moreover, Symbol associated operations resolve in a polymorphic manner before being evaluated by the NeSy engines. SymbolicAI’s engine implementation contains a prepare method to resolve and compile the engine specific representation by evaluating the Symbol-specific operations and context. For example, when utilizing GPT-4 vision as a NeSy engine backend, we compose the system and user level prompts in the prepare statement and resolve image or video related URL queries by parsing meta tags such as the <<vision: ... :>> tag using regular expressions.
@@
In SymbolicAI, operators are overloaded to facilitate transformations of Symbol objects. These operator primitives employ dynamic casting to assure type compatibility, simplifying declarations. Consequently, Symbol objects can be easily manipulated through type specific attributions or symbolically evaluated by the NeSy engine. For example, a central operation for boolean logic is measuring equality between symbols. To evaluate the equality of symbols, we primarily adhere to the type specific implementation, because we prioritize strict comparisons over probabilistic evaluations. If the evaluation was unsuccessful, we then consider semantic equality through the NeSy engine. SymbolicAI leverages decorators for compose operators and custom class methods. For more details, see supplementary materials Section C. Upon invoking an operator or method, the respective primitive function evaluates the symbol’s specific type and its respective attributes, and if further necessary, resolves a nested decorated function that then references the NeSy engine for an evaluation. Should the evaluation fail, a predefined fallback implementation executes. Absent a fallback, or if both evaluations fail, an error state is raised. The processing of an operator or custom method involves a pipeline consisting of pre- and post-processing steps, as well as constraint enforcement. Constraints cover aspects like return types, value ranges, and structural integrity (e.g. JSON formatting through grammar-based verification). In Figure 3 we give an overview of the entire prompt composition based on the user input, the Symbol object structure, and the NeSy engine evaluation pipeline.
Figure 3: We illustrate our NeSy engine pipeline and how we map user input and object instance attributes to a prompt instruction. Yellow marks user input, blue marks templating and transformations, red marks constraints, and violet the model prediction placeholder. The user input and Symbol instance attribute values are evaluated and transformed according to the pre- and post-processing steps. A custom method is constructed by using our reference implementation of decorators and passes the user input and instance attributes to the respective engine. A pre-processor manipulates the string representation of user input and Symbol instance attributes before forwarding them to the engine. For instance, in case the NeSy engine is an LLM (i.e. GPT-4), the static and dynamic context of a Symbol instance are mapped to the system and user prompts for that respective LLM using the prepare method of the engine. Other attributes, such as examples or templates, can help pre-define structures for the desired LLM output format. After the result is computed by the engine, it is passed to the post-processors and validated against the constraints.
@@
SymbolicAI augments the generative process by enabling systems to introspect and modify their behavior dynamically. We leverage LLMs to execute tasks based on both natural and formal language instructions, adhering to the specified user objectives and with innate self-referential structures. We derive subtypes from Expression and enclose their functionalities in task-specific components, which we then expose again through templating and the model-driven design of the NeSy engine. This design choice allows a system to create and utilize its own sub-process definitions, analogous to concepts discussed in Schmidhuber (2007, 2009). Concretely, we utilize generalization properties from LLMs to interpret and formulate a set of operations that incorporate self-instructions (Wang et al., 2022). Consequently, the operations hold the flexibility to adapt to the context, and derive sub-processes that self-instruct LLMs to engage in situational modeling and context-sensitive problem-solving. Ultimately, this enables the construction of hierarchical computational graphs for self-referential meta-reasoning systems without the need to explicitly training a meta-learner (Kirsch & Schmidhuber, 2022).
6 Evaluation
In this section, we conclude our paper with an empirical evaluation and potential areas of application using state-of-the-art LLMs as NeSy engine backends. For our evaluation we focus on the GPT family of models GPT-3.5 Turbo (revision 1106) and GPT-4 Turbo (revision 1106) (Brown et al., 2020) as they are the most proficient models to this date; Gemini-Pro as the best performing model available through API from Google; LlaMA 2 13B as it defines a good reference implementation for available open-source LLMs from Meta; Mistral 7B and Zephyr 7B as good baselines for revised and fine-tuned open-source contestants respectively. The selected open-source models Mistral, Zephyr, and LlaMA 2 are expected to have roughly equivalent parameter counts compared to GPT-3.5 Turbo and Gemini-Pro. All our experiments are expected to require a context size smaller or equal to $4096$ to enable the comparisons among the in-context capabilities across model architectures. For LlaMA 2 we use the chat version since it better follows instructions. Our evaluation focuses on three main parts. First, we evaluate basic capabilities of the models to verify their consistency and proficiency with in-context learning tasks. Second, we evaluate the logic capability of models by defining logical expressions in mathematical and natural language form and measure the capability of the models to translate and evaluate logical statements across domains. Lastly, we evaluate the capability of models in planning, constructing, maintaining, and executing hierarchical computational graphs. The models are instructed to extract an execution plan, which they then must schedule and execute. A significant challenge is the evaluation of the semantics of a multi-step generative process, since the naïve assessment of task succession would score all models to zero and render them as unusable. Therefore, we derive a quality measure to address these issues.
Performance Measure

### v3

@@ 3 Problem Definition
Conventional approaches employing foundation models, such as LLMs, are predominantly confined to single-step or few-step executions and primarily reliant on hand-crafted prompt instructions, often referred to as in-context learning. This restricted scope limits the utilization of different modalities, lacks verification, and exhibits limited tool proficiency. We posit that the use of NeSy engines as core computation units, realized through logic-based methodologies coupled with sub-symbolic foundation models, offers a more general, robust, and verifiable perspective. This approach has several advantages. Firstly, it enables the integration of pre-existing solutions (e.g. various classical algorithms), offloading computational complexity and bridging different modalities. Secondly, it allows sub-symbolic components to focus on decision-making (e.g. selecting the respective tool based on in-context classification). Thirdly, it provides an interpretable language-based control layer for explainable, autonomous systems. In the following section, we elaborate on the key design principles underlying SymbolicAI and how we guide the generative processes of NeSy engines. For further technical details, see Appendix Section 5.
@@ 4 Design Principles
[…]
Furthermore, we utilize language as a tool for mapping complex concepts, leveraging its inherent semantics and abstractions to describe states and properties of a problem at hand. These mappings are universal, e.g. they may be utilized to define scene descriptions, long-horizon planning, acoustic properties, emotional states, physical conditions, etc. Therefore, language serves as a comprehensive, yet abstract framework to encapsulate meanings, and refer to it as the convex hull of the knowledge of our society. Subsequently, it is common to attribute existing physical objects with abstract concepts, as exemplified by our natural tendency to link tangible objects to colors and emotions, such as blending the color ”red” with ”heart”, ”warm”, and ”passion”. This approach also anchors our work in the field of formal language theory, as we require a structured method to construct mappings from the world to language. Consequently, we use formal language structures, such as grammars, to systematically define our language-centric approach to problem-solving and the associated translation of real-world complexities into linguistic terms.
@@
In formal language theory and linguistics, languages are structured following the Chomsky hierarchy, which classifies languages by the complexity of their grammatical structure (Chomsky, 1956). This hierarchy defines four types of grammars (Type-3 to Type-0) and separates formal languages by their grammatical complexity. A grammar in this context consists of terminal and non-terminal symbols, production rules, and a designated start symbol, enabling the generation of valid strings within a language.
We define a NeSy engine as a mapping $\mathcal{V}_{\mathcal{S}^{*}}:\Omega\times N\times\mathcal{T}\rightarrow\Sigma$, where $N\subset\Sigma$ is a set of non-terminal symbols, $\mathcal{T}\subset\Sigma$ is a set of terminal symbols and $N\cap\mathcal{T}=\emptyset$, and $\mathcal{S}^{*}\in\Sigma$ is a starting symbol. We further formalize a grammar $G=(N,\mathcal{T},P,\mathcal{S}^{*})$ with production rules defined as a $P:=\mathcal{V}_{\mathcal{S}^{*}}(\omega,N,\mathcal{T})$. This grammar describes the generation of symbols through expressions $\omega$. For simplicity, we will drop the subscript of $\mathcal{V}_{\mathcal{S}^{*}}$ and use it as $\mathcal{V}$.
@@
Figure 2: Illustration for NeSy pipeline, showcasing conceptual usage of in-context learning methodologies, domain-specific language (DSL) structures, and the expression evaluations through a NeSy engine based on an LLM and constraint verification. The expression showcases the sorted insert operator $\ll$ and how the information of the symbol B is included in the symbol AC. The violet placeholder in the DSL Prompt represents an instruction, such as ”Insert the right-hand side value into the left-hand value in ascending order.” The positions below represent task-specific few-shot examples. The DSL Prompt receives the expression $\omega_{<<}$ and maps it to $\hat{\omega}_{<<}$ that can be processed by the LLM-based NeSy function $\mathcal{V}_{\mathcal{S}^{*}}$ and outputs a new symbol.
We identify LLMs as promising candidates for functioning as part of NeSy engines. In SymbolicAI, a symbol $\mathcal{S}$ is augmented with conditional instructions and types derived from DSLs, custom defined or not (e.g. HTML, SQL, etc.), tailored for directing the LLMs. The key advantage of LLMs over previous systems lies in their ability to generalize across formal languages (Wang et al., 2023a) and knowledge systems. Although there is currently no universal consensus regarding the precise classification of natural language within the Chomsky hierarchy, our approach can be understood as employing a situation-specific, context-sensitive grammar, which enables the processing of instructions and analogies with a nuanced understanding of language. The intersection between formal and natural languages becomes evident when considering how language patterns, through prompts like ”You are a helpful assistant…”, elicit structured responses, indicating a potential underlying formal mechanism at play. This observation underlines the utility of such a grammar in our framework, where it serves as an explicit schema guiding the structure of examples for in-context learning. For instance, equating ”3.1415…” with ”$\pi$” or ”August 4, 1961” with ”1961-08-04” in a given context demonstrates context-dependent interpretation of symbols. Such a system doesn’t rigidly adhere to standard grammatical rules but instead adjusts and interprets based on the context, effectively creating a situation-specific grammar, capable of forming Domain-Invariant Associations through in-context learning. We further address this in a later paragraph.
Function Composition
In SymbolicAI, we use function composition to construct complex hierarchies and behaviors from fundamental elements. Therefore, our framework enables modeling of interconnected processes, where the output of one function is used as input for another, thus creating a sequence of operations. Through function composition, we construct computational graphs, in which intermediate symbols represent the nodes or states within these graphs. Formally, function composition is denoted by $\circ$, where combining functions $f$ and $g$ yields a new function $h=g\circ f$, defined as $h(x)=g(f(x))$ For functions $f:X\rightarrow Y$ and $g:Y\rightarrow Z$, their composition results in a function mapping elements from domain $X$ to codomain $Z$ through $g(f(x))$. Although traditionally the codomain of the inner function $f$ aligns with the domain of the outer function $g$, SymbolicAI relaxes this constraint by allowing for any subset relationship between these domains and codomains, which is particularly beneficial for in-context learning. When using LLMs for NeSy production rules $\mathcal{V}$, we can derive a multi-step generative process by composing a computational graph as a sequence of zero- and few-shot function compositions:
$\mathcal{V}(\omega_{j},N,\mathcal{T})=\mathcal{V}(\omega_{j-1},\cdot)\circ\mathcal{V}(\omega_{j-2},\cdot)\circ\dots\circ\mathcal{V}(\omega_{0},\cdot),$ (1)
where $\omega_{0}$ is the initial instruction and $j$ defines the index variable for a multi-step generative process. By leveraging functional in-context learning, where zero- and few-shot examples act as dynamic elements of the function’s domain, SymbolicAI has the ability to interpret and respond to diverse input contexts. For instance, a function can classify a user request and select an appropriate interface (e.g. WolframAlpha) to process the request. The output modality may even vary based on the respective engine. This enables SymbolicAI to handle operations over multi-modal data that connects multi-step generative processes and establishes function composition as a central tenet in bridging multiple modalities and coordinating a variety of tasks.
@@
In-context learning enabled LLMs to become versatile task solvers by interpolating within the training distribution, to the extent that even potentially unseen tasks are addressable (Brown et al., 2020). We attribute this to associations formed within the input space and the capacity of Transformer architectures for defining domain-invariant feature sub-spaces. This phenomenon has parallels with few-shot learning approaches such as SubGD (Gauch et al., 2022), a method based on identifying and utilizing a low-dimensional subspace, learned from various tasks that effectively regularize the learning process. Since LLMs have been trained on different domains and tasks, which also include formulations of mathematical expressions, we posit that specific tokens, such as the equality sign, can be leveraged to associate meanings between different symbolic objects. Unlike domain-invariant representations that create invariant features across different learning tasks, our approach leverages the in-context generalization capability of LLMs to construct invariant symbolic associations that aim to preserve, manipulate and propagate situational context. We can use these properties to build operations that apply transformations on objects that are substitutes to the semantically aligned few-shot learning examples.
5 SymbolicAI Framework
In this section, we discuss the specifics of the proposed SymbolicAI framework. For more details about the framework structure, see Appendix Section C. For installation and usage of our framework, see Appendix Section D. For more technical details and code snippets, see Appendix Section E.
@@ 5 SymbolicAI Framework
Analogous to the type object in Python, the base type of SymbolicAI is a symbol represented by the base type Symbol. All other subtypes, such as Expression, represent their mathematical namesake and can be evaluated and simplified. These subtypes inherit from Symbol the base attributes, primitive operators, and helper methods.
Although SymbolicAI uses a language-centric design, modeling and manipulating every interaction into symbolic representations is not inherently efficient. Therefore, we establish mappings between symbolic and sub-symbolic representations for sensory inputs and non-discrete elements. Such mappings are typically realized through function approximation. This allows us to map between modality-to-language and language-to-modality use cases. Here, modality serves as a placeholder for various types such as text, image, video, audio, motion, etc. In turn, each Symbol object contains valued and vector-valued representations, obtained through value and embedding attributes. The latter represents a symbol’s current value, akin to embedding text and storing it as a PyTorch tensor (Paszke et al., 2019) or NumPy array (Harris et al., 2020). While for an LLM, the numerical tensors may lack inherent meaning, vector-valued representations play an important role when 1) composite symbols are combined into more complex expressions, and 2) these embedded tensors are updated through gradient-based optimization.
To enable the processing of symbols by LLMs, we assume that each Symbol object implements Python’s native string functionality, where the __str__ method returns an interpretable string representation. Therefore, we can assert that any Python object is parsable by an LLM, however, the user must ensure a meaningful representation. For more details, see Appendix Section E.
@@
Polymorphism is a central concept in programming language theory and prominently featured in SymbolicAI. Polymorphism refers to the ability of different objects to be accessed through the same interface, or of a single identifier to represent different types based on the context of execution. Providing a single interface for entities of different types allows operations to be performed in ways specific to their derived types. We therefore designed the Symbol object to contain a global context, which is composed of static and dynamic context parts, and enables this polymorphic behavior. The static context is class dependent and defined at design time. The dynamic context is runtime adaptable and can be changed to adhere to runtime specific logic and changes. Moreover, Symbol associated operations are resolved following polymorphic design before being evaluated by the NeSy engine. SymbolicAI’s engine implementation contains a prepare method to resolve and compile the engine specific representation by evaluating the Symbol-specific operations and context. For an example on polymorphic context see part a) in Figure 3.
@@
In SymbolicAI, operators are overloaded to facilitate transformations of Symbol objects. These operator primitives employ dynamic casting to ensure type compatibility. Consequently, Symbol objects can be easily manipulated through type specific attributions or symbolically evaluated by the NeSy engine. For example, a central operation for boolean logic is measuring equality between symbols. To evaluate the equality of symbols, we primarily adhere to the type specific implementation, because we prioritize strict comparisons over probabilistic evaluations. If the evaluation was unsuccessful, we then consider semantic equality through the NeSy engine. SymbolicAI leverages decorators for composing operators and custom class methods. For more details, see Appendix Section C.
Upon invoking an operator or method, the respective primitive function evaluates the symbol’s specific type and its respective attributes, and if necessary, resolves a nested decorated function that then uses the NeSy engine for evaluation. Should the evaluation fail, a predefined fallback implementation executes. Absent a fallback, or if both evaluations fail, an error state is raised. The processing of an operator or custom method involves a pipeline consisting of pre- and post-processing steps, as well as constraint enforcement. Constraints cover aspects like return types, value ranges, and structural integrity (e.g. JSON formatting through grammar-based verification). In Figure 3 b) we give an overview of the entire prompt composition based on the user input, the Symbol object structure, and in part c) the Symbol evaluation pipeline.
Figure 3: a) Illustration of polymorphic context on the example of a SQLExpression type for the add-operator. Without a polymorphic context a regular Expression evaluation concatenates two Symbol objects together. The polymorphic context in SQLExpression overwrites the base behavior such that two added SQL-expressions get semantically combined, not concatenated. b) Illustration of the translation of a Symbol object to a prompt statement to be processed by an LLM in the NeSy engine. The User Input Args can be attached with a Payload from previous executions and gets applied to the Custom Method. The user input with the polymorphic context of the Symbol Object attributes (Static Context and Dynamic Context) are translated to a prompt statement according to the schema of the Prompt Design. The fields Operation, Examples and Template mark operation description, DSL-based prompt examples and template structures respectively. These translations are processed according to PreProcessor and engine-specific formatting. c) Illustrates the evaluation pipeline from user input to output, with multiple translation processes before and after the Engine invocation. The Input gets passed to the Custom Method and reformatted according to a PreProcessor to adhere to DSL-specific structure. The engine then takes the output of the PreProcessor and composes the final prompt according to the engine-specific Prompt Design and resolves polymorphic context and auxiliary fields. The output of the Engine then can be restructured by a PostProcessor to match DSL-requirements of the desired Output and gets applied Constraints to verify the outcome.
@@
SymbolicAI augments the generative process by enabling systems to introspect and modify their behavior dynamically. We leverage LLMs to execute tasks based on both natural and formal language instructions, adhering to the specified user objectives and with innate self-referential structures. We derive subtypes from Expression and enclose their functionalities in task-specific components, which we then expose again through templating and the model-driven design of the NeSy engine. This design choice allows a system to create and utilize its own sub-process definitions, analogous to concepts discussed in Schmidhuber (2007; 2009). Concretely, we utilize generalization properties of LLMs to interpret and formulate a set of operations that incorporate self-instructions (Wang et al., 2022). Consequently, the operations hold the flexibility to adapt to the context, and derive sub-processes that self-instruct LLMs to engage in situational modeling and context-sensitive problem-solving. Ultimately, this enables the construction of hierarchical computational graphs for self-referential meta-reasoning systems without the need to explicitly training a meta-learner (Kirsch & Schmidhuber, 2022). In Figure 4 we illustrate a step-wise evaluation of a contextual computational graph, in which the NeSy engine is processing conditioned on the current execution context and producing a next symbol prediction.
6 Performance Measure